# Words Matter: Leveraging Individual Text Embeddings for Code Generation in CLIP Test-Time Adaptation

### v2

@@ 6 Conclusions
We have presented a novel approach for TTA of VLMs. In particular, CLIP-OT generates pseudo-labels for test samples using multiple, generic class templates, optimizing a label assignment problem solved with Optimal Transport. This strategy yields state-of-the-art results compared to recent works in several benchmarks, yet being a significantly more efficient solution in terms of computational complexity.
@@
Appendix A Templates Used
In the experimental setup, several predefined text templates from CLIP [11] were used to evaluate the proposed model’s adaptability and performance. These are designed to generalize across different contexts and image types, enabling a robust assessment of the model’s capabilities in handling diverse visual and textual representations. In Table 6, we provide the templates used in our experiments. Each template includes a placeholder , which is dynamically replaced with the class name during the generation of text prompts. Note that, for the sake of fairness with prior literature, the templates are the same as the ones employed in WATT [31].
@@ Appendix A Templates Used
1: a photo of a {class $k$}
2: itap of a {class $k$}
3: a bad photo of the {class $k$}
4: a origami {class $k$}
5: a photo of the large {class $k$}
6: a {class $k$} in a video game
7: art of the {class $k$}
8: a photo of the small {class $k$}
Table 6: The different templates used during the experiments.
Appendix B Additional details on the datasets
CIFAR-10.1 introduces a natural shift from CIFAR-10, whereas CIFAR-10-C and CIFAR-100-C are augmented with 15 different corruptions across 5 severity levels (each containing 10,000 images), leading to 75 corruption scenarios commonly employed in domain shift problems. These datasets are critical for evaluating how well models can generalize to real-world variations and noise that are not present in the training set. Tiny-ImageNet is a downsized version of the original ImageNet dataset, providing a more accessible challenge with 200 classes. It is a useful benchmark for lightweight models and educational purposes, making it easier to implement and experiment with domain adaptation techniques. Tiny-ImageNet-C further extends Tiny-ImageNet by incorporating various common corruptions, similar to CIFAR-10-C and CIFAR-100-C. This dataset is instrumental for testing the robustness of models against a variety of distortions in a more manageable setting. To assess the performance on class imbalanced datasets, we employ several datasets from Domainbed (PACS, OfficeHome, VisDA-C) that are often utilized for benchmarking domain adaptation algorithms, and have been recently adopted in TTA scenarios [14]. Each dataset presents unique domain shift challenges and diverse visual categories, allowing for a comprehensive evaluation of a model’s ability to generalize across different environments.
Appendix C Additional details of the ablation studies
C.1 Configuration: Baselines
In this section, we detail the configurations used to assess the impact of each main component of our approach, which we refer to Training-free OT and CLIP-OT with Average Template. Note that, in short, Training-free OT motivates the use of the pseudo-codes $\mathbf{Q}^{*}$ over the CLIP baseline, whereas CLIP-OT with Average Template includes these pseudo-codes to fine-tune the model. Furthermore, we want to highlight that the later does not fully leverage multiple individual templates, which is introduced in our proposed CLIP-OT model. A more detailed description follows. Given a batch of test images, Training-free OT (Algorithm 2) computes pseudo-codes for each text template, which are later averaged to produce a final prediction. On the other hand, CLIP-OT with Average Template (Algorithm 3) utilizes the averaged pseudo-codes (obtained by Training-free OT) to refine the visual encoder. More concretely, at each batch, the average pseudo-codes $\tilde{\mathbf{Q}}$ supervise the predictions of the test images produced by the model. These predictions are obtained by resorting to Eq. (1), where the average class embedding $\mathbf{t}_{k}$ is used, together with the visual embeddings of the test images, $\mathbf{Z}$. Differences between the pseudo-codes and the predictions are minimized via a cross-entropy loss, whose gradients are used to update the layer norm parameters of the model. Then, once the model is updated, we can do the final inference for the test images.
C.2 Extended numerical values
We further substantiate the findings presented in Figure 2 by providing detailed results in Tables 7 and 8. The tables showcase a comprehensive performance comparison of the impact of the different components of our approach, which empirically motivate our model. These results are reported for different corruption benchmarks on the CIFAR-10, CIFAR-10.1, CIFAR-10C, CIFAR-100, and CIFAR-100C datasets, respectively, using a ViT-B/32 backbone. Also, Table 9 extends the results obtained on the different domains of PACS and OfficeHome, whose average results were reported in the main paper. We can observe that only in one out of the different datasets and domains, CLIP-OT underperforms previous SoTA, with a marginal performance gap, i.e., -0.19%.
@@ C.2 Extended numerical values
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{s}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do
4:       for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
@@
6:             $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:             Compute pseudo-codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
@@
9:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict: $\mathbf{p}_{i}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{q}_{im}$.
@@
Algorithm 3 CLIP-OT with Average Template.
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{s}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do // 1 - Adaptation
4:       for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
@@
6:             $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:             Compute pseudo-codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
8:       end for // Step 2: Compute average pseudo-codes and refine the encoder.
9:       $\tilde{\mathbf{Q}}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{Q}^{*}_{m}$ // average pseudo-codes.
10:       $\mathbf{t}_{k}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{t}_{km}$ // average class text embeddings over all templates.
11:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
@@
14:       $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features (with $\boldsymbol{\theta}_{\text{LN}}^{(m)}$).
15:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
@@
Dataset CLIP Training-Free OT Avg. Template CLIP-OT
@@
Table 7: Numerical Analysis for CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets as shown in Figure 2.
Corruption CLIP Training-Free OT Avg. Template CLIP-OT
@@
Dataset Domain CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
PACS Art 96.34 96.65 $\pm$0.05 95.52 $\pm$0.20 96.57 $\pm$0.09 96.31 $\pm$0.00 96.39 $\pm$0.00 96.54 $\pm$0.05 +0.15 $\blacktriangle$
Cartoon 96.08 96.22 $\pm$0.05 94.77 $\pm$0.20 96.00 $\pm$0.02 96.52 $\pm$0.02 96.62 $\pm$0.02 98.34 $\pm$0.13 +1.72 $\blacktriangle$
Photo 99.34 99.40 $\pm$0.00 99.42 $\pm$0.06 99.28 $\pm$0.00 99.48 $\pm$0.03 99.52 $\pm$0.00 99.86 $\pm$0.03 +0.34 $\blacktriangle$
Sketch 82.85 82.96 $\pm$0.12 83.22 $\pm$0.14 83.93 $\pm$0.14 86.92 $\pm$0.04 86.65 $\pm$0.12 89.01 $\pm$0.29 +2.36 $\blacktriangle$
Mean 93.65 93.81 93.23 93.95 94.81 94.80 95.94 +1.14 $\blacktriangle$
OfficeHome Art 73.75 74.03 $\pm$0.27 75.76 $\pm$0.27 73.84 $\pm$0.20 75.65 $\pm$0.27 75.76 $\pm$0.39 79.39 ±$\pm 0.31$ +3.63 $\blacktriangle$
Clipart 63.33 63.42 $\pm$0.04 63.08 $\pm$0.31 63.54 $\pm$0.06 66.23 $\pm$0.13 65.77 $\pm$0.11 66.21 $\pm$0.17 +0.44 $\blacktriangle$
Product 85.32 85.51 $\pm$0.08 84.07 $\pm$0.28 85.23 $\pm$0.16 85.41 $\pm$0.09 85.41 $\pm$0.01 86.81 $\pm$0.03 +1.40 $\blacktriangle$
[…]
Mean 77.53 77.68 77.20 77.56 78.88 78.83 80.15 +1.32 $\blacktriangle$
VisDA-C 3D (trainset) 84.43 84.86 $\pm$0.01 79.35 $\pm$0.04 85.09 $\pm$0.01 85.42 $\pm$0.03 85.36 $\pm$0.01 90.73 $\pm$0.65 +5.37 $\blacktriangle$
YT (valset) 84.45 84.68 $\pm$0.01 83.57 $\pm$0.04 84.40 $\pm$0.01 84.57 $\pm$0.00 84.69 $\pm$0.01 85.44 $\pm$0.65 +0.75 $\blacktriangle$
Mean 84.44 84.77 81.46 84.75 85.00 85.03 88.08 +3.05 $\blacktriangle$
[…]
Appendix D Comparison to OT-VP
The very recent OT-VP [48] has presented a solution integrating optimal transport for test-time adaptation. Nevertheless, it presents several fundamental differences with our work. First, OT-VP involves learning a universal visual prompt for the target domain, for which an optimal transport distance is optimized. Thus, optimal transport is used for a different task. Second, OT-VP is tailored to only visual models, not being capable of leveraging the valuable information found on the text modality. And third, the adaptation scenario strongly differs from our setting. In particular, OT-VP first fine-tunes the pre-trained model (pre-trained on ImageNet-1k) to a subdomain (e.g., PACS) with labeled data, to later adapt at test-time to the other subdomains (e.g., OfficeHome). While this strategy can be done in a single-source setting, OT-VP also evaluates the performance when multiple domains are used as the source (e.g., supervised adaptation on PACS), and only one left out for testing, referred to as Multi-Source in Table 10. In contrast, we follow the protocol of recent literature of CLIP test-time adaptation, where CLIP is directly exposed to the unsupervised test data points, without intermediate adaptation steps. To expand the extent of our empirical validation, we compare our approach to OT-VP [48] whose results are reported in Table 10. To ensure a rigorous comparison, we use the same visual encoder as the backbone used in OT-VP and other prior models, i.e., ViT-B/16.
Method PACS OfficeHome Avg.
First Setting : Single Source
[…]
Second Setting : Multi Source
[…]
Our setting: No specific source
CLIP-OT(Ours) 96.9+23.4/+9.2 83.4+15.3/+8.3 86.8+16.8/+5.6
Table 10: Performance comparison to [48] across PACS and OfficeHome datasets. Best results highlighted in bold, whereas value indicates the difference wrt OT-VP in both single (first value) and multi (second value) source scenarios.
Appendix E Differences wrt PLOT [6]
Note that despite the apparent similarities, our work presents fundamental differences w.r.t. Prompt Learning with Optimal Transport (PLOT) [6], which prevents from adding the later in the empirical evaluation. a) Target task. PLOT tackles few-shot adaptation, requiring labeled data to adapt no novel tasks, whereas our method studies test-time adaptation, where labels or test samples are not available. Thus PLOT cannot be compared in the TTA setting. b) Adaptation. Our approach updates the layer norm parameters of CLIP visual encoder, whereas PLOT optimizes the input text prompts (i.e., it falls into the Prompt Learning category). c) Use of OT. PLOT leverages OT between learnable prompts and a set of local visual features, whose transport plan is directly used as predicted probabilities. In contrast, we compute the transportation cost between the whole image features and fixed text embeddings, which is later used as a pseudo-supervision to distill knowledge during test-time adaptation.
Appendix F Additional Details on Epsilon ($\epsilon$)
@@
Table 11: Mean Accuracy for CIFAR-10 and CIFAR-100 under Different Corruptions and Epsilon Values

### v3

@@ 6 Conclusions
We have presented Semantic Anchor Transport (SAT), a novel approach for the test-time adaptation of vision-language models. SAT reformulates TTA as a principled cross-modal alignment problem. It generates robust, batch-aware pseudo-labels by aligning visual embeddings to fixed text-based semantic anchors using Optimal Transport. This global assignment strategy fundamentally mitigates the error accumulation demonstrated by other TTA methods. Furthermore, SAT employs a sophisticated multi-template distillation strategy to harness diverse textual clues, enhancing robustness without incurring significant computational overhead. Extensive experiments demonstrate that SAT achieves state-of-the-art performance on TTA across multiple visual domain shift benchmarks and multi-modal backbones.
[…]
@@
A Additional datasets details.
Datasets. CIFAR-10.1 introduces a natural shift from CIFAR-10, whereas CIFAR-10-C and CIFAR-100-C are augmented with 15 different corruptions across 5 severity levels (each containing 10,000 images), leading to 75 corruption scenarios commonly employed in domain shift problems. These datasets are critical for evaluating how well models can generalize to real-world variations and noise that are not present in the training set. Similarly, ImageNet-C provides these corruptions for the 1000-class ImageNet validation set. These datasets are critical for evaluating how well models can generalize to real-world variations and noise. Tiny-ImageNet is a downsized version of the original ImageNet dataset, providing a more accessible challenge with 200 classes. Tiny-ImageNet-C further extends Tiny-ImageNet by incorporating various common corruptions. This dataset is instrumental for testing robustness against a variety of distortions. To assess the performance on class imbalanced datasets, we employ several datasets from Domainbed (PACS, OfficeHome, VisDA-C) that are often utilized for benchmarking domain adaptation algorithms, and have been recently adopted in TTA scenarios [hakim2024clipartt]. Each dataset presents unique domain shift challenges and diverse visual categories, allowing for a comprehensive evaluation of a model’s ability to generalize across different environments.
CLIP’s text templates. In the experimental setup, several predefined text templates from CLIP [clipad] were used to evaluate the proposed model’s adaptability and performance. These are designed to generalize across different contexts and image types, enabling a robust assessment of the model’s capabilities in handling diverse visual and textual representations. In Table 5, we provide the templates used in our experiments. Each template includes a placeholder , which is dynamically replaced with the class name during the generation of text prompts. Note that, for the sake of fairness with prior literature, the templates are the same as the ones employed in WATT [osowiechi2024watt].
Table 5: The different templates used during the experiments.
@@
1: a photo of a {class kk}
2: itap of a {class kk}
3: a bad photo of the {class kk}
4: a origami {class kk}
5: a photo of the large {class kk}
6: a {class kk} in a video game
7: art of the {class kk}
8: a photo of the small {class kk}
B Additional experimental details.
B.1 Configuration: Baselines
In this section, we detail the configurations used to assess the impact of each main component of our approach, which we refer to Training-free OT and SAT with Average Template. Note that, in short, Training-free OT motivates the use of the soft assignments $\mathbf{Q}^{*}$ over the CLIP baseline, whereas SAT with Average Template includes these assignments to fine-tune the model. Furthermore, we want to highlight that the latter does not fully leverage multiple individual templates, which is introduced in our proposed method. A more detailed description follows. Given a batch of test images, Training-free OT (Algorithm 2) computes soft assignments for each text template, which are later averaged to produce a final prediction. On the other hand, SAT with Average Template (Algorithm 3) utilizes the averaged assignments (obtained by Training-free OT) to refine the visual encoder. More concretely, at each batch, the average assignments $\tilde{\mathbf{Q}}$ supervise the predictions of the test images produced by the model. These predictions are obtained by resorting to Eq. (1), where the average class embedding $\mathbf{T}$ is used, together with the visual embeddings of the test images, $\mathbf{Z}$. Differences between the assignments and the predictions are minimized via a cross-entropy loss, whose gradients are used to update the layer norm parameters of the model. Then, once the model is updated, we can do the final inference for the test images.
@@ B.1 Configuration: Baselines
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{T}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
[…]
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{T}}$ do
4:   for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Align - Compute soft assignments $\mathbf{Q}^{*}_{m}$.
@@
6:     $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{T}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:     Compute soft assignments $\mathbf{Q}^{*}_{m}$ // Eq. (9).
@@
9:   $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{T}}]$ // predict: $\mathbf{p}_{i}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{q}_{im}$.
@@
Algorithm 3 SAT with Average Template.
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{T}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
[…]
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{T}}$ do // 1 - Adaptation
4:   for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Align - Compute soft assignments $\mathbf{Q}^{*}_{m}$.
@@
6:     $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{T}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:     Compute soft assignments $\mathbf{Q}^{*}_{m}$ // Eq. (9).
8:   end for// Step 2: Compute average assignments and refine the encoder.
9:   $\tilde{\mathbf{Q}}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{Q}^{*}_{m}$ // average assignments.
10:   $\mathbf{T}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{T}_{m}$ // average class text embeddings.
11:   $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{T}}]$ // predict with $\mathbf{T}$, Eq. (1).
@@
14:   $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{T}}]$ // visual features (with $\boldsymbol{\theta}_{\text{LN}}^{(m)}$).
15:   $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{T}}]$ // predict with $\mathbf{T}$, Eq. (1).
@@
B.2 Extended numerical values
We further substantiate the findings presented in Figure 3 by providing detailed results in Tables 6 and 7. The tables showcase a comprehensive performance comparison of the impact of the different components of our approach, which empirically motivate our model. These results are reported for different corruption benchmarks on the CIFAR-10, CIFAR-10.1, CIFAR-10C, CIFAR-100, and CIFAR-100C datasets, respectively, using a ViT-B/32 backbone.
Table 6: Numerical Analysis for CIFAR-10, CIFAR-10.1 and CIFAR-10-C as shown in Figure 3.
Dataset CLIP Training-Free OT Avg. Template SAT
@@ B.2 Extended numerical values
Corruption CLIP Training-Free OT Avg. Template SAT
@@
B.3 Additional Details on Epsilon ($\epsilon$)
@@
Figure 5: Ablation on epsilon values across multiple datasets.
Table 8: Accuracy on CIFAR-10 and CIFAR-100 under different corruptions and epsilon values.